OpenStudy (anonymous):

Simplify this expression. 3^5 ÷ 3^3 a.) 27 b.) 9 c.) 6 d.) 3

OpenStudy (anonymous):

Well what is 3^5?

OpenStudy (anonymous):

Also, what is 3^3?

OpenStudy (anonymous):

don't you dare joylin its my turn

OpenStudy (anonymous):

15 9

OpenStudy (anonymous):

Powers mean that you multiply the base by however many times the power states.

OpenStudy (anonymous):

This isn't multiplication. It's exponents. xD Last one, KD.

OpenStudy (anonymous):

So 3^3=(3)(3)(3)

OpenStudy (anonymous):

I know i was wondering if I subtracted the exponents???

OpenStudy (anonymous):

And 3^5=(3)(3)(3)(3)(3)

OpenStudy (anonymous):

:o how dare you joylin

OpenStudy (anonymous):

Oh oops I thought you said 3*5! Opps!

OpenStudy (anonymous):

So now it's...